OpenStudy (anonymous):

Hey guys l need help on this x^2+3x+2 So what I'm suppose to be doing is to graph quadratic functions.... So can you please help? (: I would appreciate it (:

OpenStudy (anonymous):

what multiplies to make 2 and adds to make 3?

OpenStudy (anonymous):

or what are the factors of 2?

OpenStudy (anonymous):

first thing you need to do is solve for when y= zero to find x intercepts

OpenStudy (anonymous):

you do this by factoring which is why I asked what adds to make 3 and multiplies to make 2

OpenStudy (anonymous):

next you need to to find y intercept by setting x= 0 and you mark that on the graph along with the x intercepts

OpenStudy (anonymous):

then you work your way towards the y intercept from either side by filling in different values of x do you know what I mean?

OpenStudy (anonymous):

But is the Vertex (-1.5, 0.5)?

OpenStudy (anonymous):

1.find the x and y intercepts y = 2 when u substitute 0 on x (0,2) x= -2 or x = -1 factorise (-2,0) and (-1,0) 2.find the turning point (quadratic eqn will have one turning point) differentiate ur function ,equate to zero and solve for x f'(x) = 2x +3 f'(x) = 0 2x + 3 = 0 x =-3/2 find y y = (-3/2)^2 + 3(-3/2) then u get the coordinates of ur turning point 3.check if ur graph is a concave up or concave down it is concave up if ur function is increasing (f' > 0) and concave down if ur function is decreasing (f'<0) now u can plot ur graph using intercepts ,and turning point hope i was of help

OpenStudy (anonymous):

lol bro she or he is not about to differentiate anything just saying this clearly is not from a calculus class or is it?

OpenStudy (anonymous):

oh sorry my bad I misinformed you ignore the y-int part but for a quadratic equation should be in the middle of the two x-int if you know how to differentiate this will give you the vertex as well. by finding when f'(x) = 0 and then subbing that x value back into the original equation like bekesto said.

OpenStudy (anonymous):

wait can please explain again?